OpenStudy (anonymous):

Can someone help me solve this problem? An Initial investment of $1000 is appreciated for 8 years in an account that earns 9% interest, compounded annually. Find the amount of money in the account the end of the period.

OpenStudy (amistre64):

to compund is to re evaluate the amount with the added interest

OpenStudy (amistre64):

compound even

OpenStudy (anonymous):

\[1000*(1.09)^8\]

OpenStudy (anonymous):

Initially you have 1000 1 year later you have 9% more 1000*1.09 2 years later: (1000*1.09)*1.09 = 1000*1.09^2 And so on.

OpenStudy (amistre64):

A(0) = 1000 A(1) = 1000 + 1000(.09) A(2) = 1000 + (1000 + 1000(.09)) (.09) = 1000 + 1000(.09) + 1000(.09)(.09) = 1000 + 1000(.09) + 1000(.09^2) = 1000 + 1000(.09 + .09^2) = 1000(1 + (.09 + .09^2)) A(3) = 1000(1 + (.09 + .09^2 + .09^3)) etc

OpenStudy (curry):

nice amis

OpenStudy (amistre64):

1.00 Sn = (.09 + .09^2 + .09^3 + ... + .09^(n)) -.09 Sn = ( - .09^2 - .09^3 - ... - .09^(n) - .09^(n+1)) ---------------------------------------------------- 1-.09 Sn = (.09 - .09^(n+1)) ; divide off the (1-.09) Sn = (.09 - .09^(n+1))/(1-.09) = .09 (1-.09^n)/(1-.09) when n=8; we get .09(1-.09^8)/(1-.09) =abt. 0.0989 A(8) = 1000(1+.0989) = 1098.90 hmmm .... i missed it along the way someplace :) the right answer is given by the fiddler: A(8) = 1000(1+.09)^8 = 1992.56

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Compound_interest